# X-TIME: An in-memory engine for accelerating machine learning on tabular data with CAMs

### v3

@@ V-B Comparison with state-of-the-art
Figure 10: Comparison of (a) latency and (b) throughput of this work, GPU, FPGA, and SRAM based accelerator from Fig. 8 (a) Throughput as a function of $N_{trees}$ and $D$.
@@
We benchmarked our proposed architecture on a wide range of models (corresponding to X-TIME 8bit of Fig. 9) and dataset types (Table II), and compared latency and throughput (see Fig. 10(a,b)) to GPU, FPGA and a digital accelerator based on small SRAMs and similar to Booster [5]. When possible, we applied input batching and replicated trees to increase throughput. As shown, our architecture produces drastically reduced latency, frequently to $\sim$100ns compared to GPU at 10 $\mu$s to ms, while also demonstrating significantly improved throughput by 10-120$\times$ across most datasets. Our architecture shines for large binary classification or regression models with a peak improvement in latency by 9740×9740$\times$9740 × at 119$\times$ higher throughput for the Churn modeling dataset inference[33] compared to GPU. As tree search is parallel in each analog CAM array of each core, as well as the in-network reduction, latency, and throughput are constant with $N_{trees}$ and $N_{leaves}$ with a delay solely dependant on $N_{feat}$ and $N_{classes}$. In contrast, on GPU there is a linear dependence on $N_{trees}$ and $D$, which limits the performance for large models. In the FPGA implementation, namely FAXID, there is a linear dependence on $N_{trees}$. The SRAM-based accelerator has a linear dependence on $D$ (since multiple memory accesses are required) and $N_{trees}$ when more than a tree needs to be mapped in a core. As an example, we studied the throughput as a function of $N_{trees}$ and $D$ for GPU, FPGA, SRAM-based accelerator, and X-TIME as shown in Fig. 10(c), for a multiclass classification dataset [35], demonstrating the predicted trend. We note that the SRAM-based accelerator still suffers from load imbalance between trees, similar to GPU, in which trees with different depths have different inference times, which results in the need for synchronization before the reduction operation. The high throughput of X-TIME can produce an extremely low energy per decision, down to 0.3 nJ/Dec, highlighting the potential for energy-efficient applications.
VI Conclusions
@@
In this work, we presented a novel architecture for accelerated tree-based ML inference with in-memory computing based on a novel increased precision analog non-volatile CAMs. Our approach maps decision trees to analog CAM ranges, which compute directly the matched results in parallel given an input feature vector. Leaf values are stored in a local SRAM and accumulated at the core level when multiple trees are mapped in the same core. Outside the core, an H-tree NoC connects 4096 cores and a co-processor in a single chip. An in-network computing structure accumulates leaf values as data flows to limit data movement. The on-chip network is re-configurable to map different models (regression, multi-class classification, etc) and inference techniques, such as input batching. Our results show significant improvement in latency – by almost four orders of magnitude, compared with an NVIDIA V100 GPU, at a throughput up to 120$\sim$ 120$\times$ 120 × higher across a wide range of dataset types and model topologies.

### v4

@@ 5.1 Comparison with state-of-the-art
Figure 7: Comparison of (a) latency and (b) throughput of this work, GPU, FPGA, and fully digital ASIC accelerator from Fig. LABEL:supp-si:fig:digital_accelerator.
@@
We benchmarked X-TIME using our open-sourced cycle-approximate simulator [^1] https://github.com/HewlettPackard/X-TIME on a wide range of models and dataset types (Table 1), and compared latency and throughput (see Fig. 7(a,b)) to GPU, specifically NVIDIA V100, FPGA [9] and a digital accelerator based on small SRAMs and similar to Booster [6] (Fig. LABEL:supp-si:fig:digital_accelerator). More details on the methodology can be found in Supplementary Information LABEL:supp-si:section:soa, LABEL:supp-si:section:methods and Table LABEL:supp-si:tab:benchmark. When possible, we applied input batching and replicated trees to increase throughput (Fig. LABEL:supp-si:fig:batching). As shown, our architecture produces drastically reduced latency, frequently to $\sim$100ns compared to GPU at 10 $\mu$s to ms, while also demonstrating significantly improved throughput by 10-120$\times$ across most datasets. Our architecture shines for large binary classification or regression models with a peak improvement in latency by 9740×9740$\times$9740 × at 119$\times$ higher throughput for the Churn modeling dataset inference [32] compared to GPU. As tree search is parallel in each analog CAM array of each core, as well as the in-network reduction, latency, and throughput are constant with $N_{trees}$ and $N_{leaves}$ with a delay solely dependant on $N_{feat}$ and $N_{classes}$. In contrast, on GPU there is a linear dependence on $N_{trees}$ and $D$, which limits the performance for large models. In the FPGA implementation, namely FAXID, there is a linear dependence on $N_{trees}$. The SRAM-based accelerator has a linear dependence on $D$ (since multiple memory accesses are required) and $N_{trees}$ when more than a tree needs to be mapped in a core.
Figure 8: Throughput as a function of number of trees $N_{trees}$ and depth $D$ for various accelerators. Note that in the case of X-TIME and the FPGA implementation, there is no dependence on the depth, removing the load imbalance issue, while only X-TIME shows no dependence on the number of trees.
As an example, we studied the throughput as a function of $N_{trees}$ and $D$ for GPU, FPGA, SRAM-based accelerator, and X-TIME as shown in Fig. 8, for a multiclass classification dataset [34], demonstrating the predicted trend. We note that the SRAM-based accelerator still suffers from load imbalance between trees, similar to GPU, in which trees with different depths have different inference times, which results in the need for synchronization before the reduction operation. As an energy comparison with GPU, we considered a worst-case scenario. Fig. LABEL:supp-si:fig:energy shows the energy saving in comparison by considering the peak power consumption (19 W) and idle power consumption (25 W) in the case of X-TIME and GPU, respectively. Even by assuming a full chip utilization for X-TIME and no energy consumption for the tree inference on GPU, X-TIME is up to 157$\times$ more energy efficient.
5.2 Sensitivity to noise
Figure 9: Accuracy and RMSE for different noise configurations
Analog hardware is subject to noise and non-idealities. However, tree-based ML ensembles are resilient to variation thanks to the averaging mechanism happening between trees of multiple boosting rounds. Fig. 9 shows the average accuracy of the models for different noise profiles, namely no noise (ideal), only ReRAM or DAC noise, and both noise sources, over 100 inference experiments. A printout of the data can also be found in Table  LABEL:supp-si:tab:noise. We characterized the noise of TaOx-based ReRAM devices at various conductance values ranging from $<0.1$ nS to 500 $\mu$S [27], by reading them 1000 times and extracting mean and standard deviation. We measured a maximum relative standard deviation of noise of $\sigma_{G}/G=0.1$. While DAC designs can reach higher precision, only 4-bits are needed for X-TIME inference. Nevertheless, we considered a pessimistic standard deviation of noise for the DAC $\sigma_{DAC}=50mV$. Results indicate no perturbation in the accuracy and RMSE when either ReRAM or DAC noise is applied, thanks to our hardware-aware training and inherent averaging mechanisms of ensemble models. In the case of regression and with both noise sources applied there is an RMSE perturbation, but still reduced compared to the 4-bit operation (Fig. 4).
Intuitively, with the thresholding mechanism of Analog CAM, it is possible to program conductances with a separation such that the probability of exceeding the threshold is relatively low, leading to good noise resistance performance.
6 Conclusion
@@
In this work, we presented a novel architecture for accelerated tree-based ML inference with in-memory computing based on a novel increased precision analog non-volatile CAMs. Analog CAMs enable novel applications of in-memory computing beyond matrix multiplies. Moreover, the required peripherals are significantly reduced compared to crossbar-based computing primitives. Our approach maps decision trees to analog CAM ranges, which directly compute the matched results in parallel, given an input feature vector. Leaf values are stored in a local SRAM and accumulated at the core level when multiple trees are mapped in the same core. Outside the core, an H-tree NoC connects 4096 cores and a co-processor in a single chip. An in-network computing structure accumulates leaf values as data flows to limit data movement. The on-chip network is re-configurable to map different models (regression, multi-class classification, etc) and inference techniques, such as input batching. An HW-aware training routine ensures to gets state of the art accuracy by leveraging quantization and optimizing the number of trees and maximum depth. Our results show significant improvement in latency – by almost four orders of magnitude, compared with an NVIDIA V100 GPU, at a throughput up to 120$\sim$ 120$\times$ 120 × higher across a wide range of dataset types and model topologies. Further, the worst case scenario showed up to 150×150$\times$150 × improvement in energy efficiency compared to GPUs with our approach.